# Learning to Play Piano in the Real World

### v1

@@ Results.
Setup. During this experiment, we trained 11 different configurations with an increasing intensity of domain randomization. To quantify the intensity of domain randomization, we introduce the parameter $C_{dr}\in[0,1]$, with $C_{dr}=0$ representing no domain randomization and $C_{dr}=1$ representing the maximum amount of domain randomization. The models are then evaluated using real world execution. The figure shows the average and standard deviation of multiple runs. To reduce the tear on the robot and since the most interesting runs are the best performing runs, we decided to reduce the number of runs for higher and lower intensities of domain randomization. We calculate the average of at least 3 runs for each configuration. Further, for $C_{dr}\in[0.2,0.8]$ we evaluated at least 5 runs and for $C_{dr}\in[0.3,0.7]$ we evaluated 7 runs. The models are distributed across 4 different seeds.
Figure 6: The diagram shows the effect of domain randomization on the performance in simulation. The more domain randomization is used, the more demanding the environment is. Since the number of training iterations is constant, this leads to a drop in performance.
Figure 7: This diagram shows the effect of domain randomization on the model performance in the real world using real world execution. The precision has a progression similar as in Fig. 6, which indicates a constant Sim2Real gap. However, the recall improves significantly until a maximum at $0.6$ where it starts dropping again. This is because the model gets more robust when more domain randomization is applied. Unfortunately, with too much domain randomization is the model not able to play piano successfully - neither in simulation nor in the real world.
Results. As visualized in Fig. 6, the performance in simulation decreases the more domain randomization is applied during training. This is because the environment behaves differently for every iteration when domain randomization is used, which requires a more robust model. When the number of iterations during the training stays the same, but the environment gets more demanding, a performance decrease is to be expected.
Further, we observe that the recall in simulation is higher than the precision. A possible explanation therefore is that our physics framework is based on soft body physics. This causes a tendency to press multiple keys at once, while only one key is intended to be pressed. This behavior ultimately reduces the precision while maintaining a high recall score.
The real world performance, on the other side, is displayed in Fig. 7. In contrast to the observations in simulation is the recall significantly lower than the precision score in the real world. This is because in the real world it is hard to press multiple keys with one finger. The plastic surfaces of the keys and fingers are slippery on each other, which causes the finger to slide onto one key. Once a key is pressed, the robot needs to lift the finger to press another key because the body of the adjacent keys acts as a border. This makes it harder to press the targeted key, since the agent might slide to the wrong side, reducing the recall score significantly.
The progression of the precision in the real world (Fig. 7) is similar to the progression of the precision in the simulation (Fig. 6). However, the real world precision is decreased by $\approx 0.2$ for all intensities.
The recall on the other side has a very different progression. Without domain randomization, the recall is at $\approx 0.44$. Increasing the intensity of domain randomization leads to a reduced recall score of $\approx 0.35$ at $C_{dr}=0.2$. From there, the recall increases significantly until the maximum of $\approx 0.6$ is reached at $C_{dr}=0.6$, which also leads to the highest F1 score of $\approx 0.52$. For $C_{dr}>0.6$ the recall drops rapidly.
A possible explanation for the similar progression of the precision is that an unintentionally pressed key can be corrected by just lifting the finger. This skill is simple to learn even without applying domain randomization. The recall on the other side, significantly benefits from the domain randomization. To correct from a mistake, the agent needs to move the finger to a different key. This requires more precise controls compared to just moving the finger up, which can be achieved by making the model more robust with domain randomization.
However, the decreased recall performance at $C_{dr}=0.2$ is harder to explain. Our understanding is, that the agent seldom ends up in those wrong states without domain randomization. Therefore, the agent is also not able to learn how to recover from those states.
For $C_{dr}>0.6$ is a drop in performance expected since the performance in simulation drops as well.
@@
Conclusion. The best performance can be achieved with moderate domain randomization. Too little domain randomization leads to worse performance since the model is not able to recover successfully after the environment behaved in a different way compared to the training. Too much domain randomization also worsens the performance since the model is not able to learn successfully during training. As a consequence of these results, we recommend researchers to carefully tune the intensity of domain randomization, since it significantly impacts the real world performance.
@@ V Limitations and Future Work
In this work, we share our findings on how to teach a robot to play simple songs on a real piano. We demonstrated a proof of concept that utilized Sim2Real and a hybrid execution mode to play piano in the real world. However, our system possesses several limitations that could become the target of future research:
@@
Fingers friction. The current 3D-printed fingertip used in our experiments is made from PLA which is very stiff. This reduces the contact area of the fingers, which reduces the friction, which leads to sliding on the keys while playing a song. This makes playing piano more demanding since the agent needs to balance the forces more precisely. Therefore, using fingertips made of softer materials, like TPU, silicone, or rubber, might increase the performance.
Bimanual playing. More complex piano songs require the use of two hands coordinating and playing at the same time. Our hardware setup is currently limited to a single hand, and this resulted in the use of relatively simple songs in our experimental setting. In the future, we hope to extend to the bimanual case and investigate how to learn more complex piano movements – thus being able to increase the overall difficulty of the played songs.
@@ VI Conclusion
In this work, we showcase the first proof-of-concept of how to learn to play piano with a real world robotic hand. Specifically, we employ a Sim2Real approach where we first train a policy in simulation, and following we deploy it on the real world dexterous hand. Experimental results show that the learned policy can play several different pieces with satisfying musical results. Nonetheless, there is still a long way before we can achieve in robots the same level of proficiency and elegance of human pianists. Beyond the artistic raison d’être, we believe that playing piano is a relevant and challenging benchmark for dexterous contact-reach manipulation, and advances in this task can be more widely relevant for the whole manipulation community.
@@ VI-A The Reward Formulation
The $r_{\text{energy}}$ reward penalizes fast movements, that need a lot of force. Further, it avoids oscillations of fingers that don’t lead to higher rewards. The energy reward also leads to a policy that behaves more fluently, which makes it more human-like.
@@
$r_{\text{hand\_position}}=\frac{1}{|P_{\text{targets}}|}*\sum_{p_{\text{key}}\in P_{\text{targets}}}g(||p_{\text{key}}-p_{\text{palm}}||_{2})$
The distances are bounded between 00 and $1$ using the tolerance function of the dm_control library [14], notated as $g$. Further is $P_{\text{targets}}$, the set of the cartesian positions of all targeted keys, and $p_{\text{palm}}$ represents the cartesian position of the palm.
@@
For $k_{\text{target}}>0$ we divide the keypress reward into tree cases

### v2

@@ IV Experimental Results
Setup. During this experiment, we trained 11 different configurations with an increasing intensity of DR. To quantify the intensity of DR, we introduce the parameter $C_{dr}\in[0,1]$, with $C_{dr}=0$ representing no DR and $C_{dr}=1$ representing the maximum amount of DR. The models are then evaluated using real world execution. The figure shows the average and standard deviation of multiple runs. To reduce the tear on the robot and since the most interesting runs are the best performing runs, we decided to reduce the number of runs for higher and lower intensities of DR. We calculate the average of at least 3 runs for each configuration. Further, for $C_{dr}\in[0.2,0.8]$ we evaluated at least 5 runs and for $C_{dr}\in[0.3,0.7]$ we evaluated 7 runs. The models are evenly distributed across 4 different seeds.
(a) The diagram shows the effect of DR on the performance in simulation.
(b) This diagram shows the effect of DR on the model performance in the real world using real world execution.
Figure 5: The more DR is applied, the more demanding the simulation environment becomes. This leads to a drop in performance, as observed in Fig. 5(a). However, DR leads to a more robust model, which can be observed in Fig. 5(b). With too much DR, the model is not able to play piano successfully, neither in simulation nor in the real world.
Results. As visualized in Fig. 5(a), the performance in simulation decreases the more DR is applied during training. This is because the environment behaves differently for every iteration when DR is used, which requires a more robust model. When the number of iterations during the training stays the same, but the environment gets more demanding, a performance decrease is to be expected. Further, we observe that the recall in the simulation is higher than the precision. A possible explanation is that our physics framework is based on soft-body physics. This introduces a tendency to press two neighboring keys at once during training, while only one key is intended to be pressed. This effect reduces the precision while maintaining a high recall score. The real world performance, on the other hand, is displayed in Fig. 5(b). In contrast to the observations in simulation, the recall is significantly lower than the precision score in the real world. This is because, in the real world, it’s difficult to press multiple keys using just one finger. The plastic surfaces of the keys and fingers are slippery on each other, which causes the finger to slide onto one key. Once a key is pressed, the robot needs to lift the finger to press another key because the body of the adjacent keys acts as a border. This makes it harder to press the targeted key, since the agent might slide to the wrong side, while in simulation, the robot would have pressed both keys. This ultimately reduces the recall score in the real world. The progression of the precision in the real world (Fig. 5(b)) is similar to the progression of the precision in the simulation (Fig. 5(a)). However, the real world precision is decreased by $\approx 0.2$ for all intensities. The recall on the other side has a very different progression compared to the simulation. Without DR, the recall is at $\approx 0.44$. Increasing the intensity of DR leads to a reduced recall score of $\approx 0.35$ at $C_{dr}=0.2$. From there, the recall increases significantly until the maximum of $\approx 0.6$ is reached at $C_{dr}=0.6$, which also leads to the highest F1 score of $\approx 0.52$. For $C_{dr}>0.6$ the recall drops rapidly. A possible explanation for the similar progression of the precision is that an unintentionally pressed key can be corrected by just lifting the finger. This skill is simple to learn even without applying DR. The recall on the other side, significantly benefits from the DR. To correct a wrong keypress, the agent needs to move the finger to a different key. This requires more precise controls compared to just moving the finger up, which can be achieved by making the model more robust with DR. However, the decreased recall performance at $C_{dr}=0.2$ is harder to explain. Our understanding is that the agent seldom ends up in those wrong states without DR. Therefore, the agent is also not able to learn how to recover from those states. For $C_{dr}>0.6$, a drop in performance is expected since the performance in simulation drops as well.
@@
Conclusion. The best performance can be achieved with moderate DR. Too little DR leads to worse performance since the model is not able to recover successfully after the environment behaved in a different way compared to the training - the model is not robust enough. Too much DR on the other side worsens the performance since the model is not able to learn successfully during training. As a consequence of these results, we recommend researchers to carefully tune the intensity of DR, since it significantly impacts the real world performance.
@@ V Limitations and Future Work
In this work, we share our findings on how to teach a robot to play simple songs on a real piano. We demonstrated a proof of concept that utilized Sim2Real and a hybrid execution mode to play piano in the real world. However, our system possesses several limitations that could be investigated in future research:
@@
Sliding Fingers. The current 3D-printed fingertip used in our experiments is made from PLA which is very stiff. This reduces the contact area of the fingers, which reduces the friction, which leads to sliding on the keys while playing a song. This makes playing the piano more demanding since the agent needs to balance the forces more precisely. Therefore, using fingertips made of softer materials, like TPU, silicone, or rubber, might increase the performance.
Bimanual Playing. More complex piano songs require the use of two hands. This also introduces the challenge of coordinating twice the number of joints and enables more complex fingering techniques. However, currently our hardware setup is limited to a single hand. In the future, we hope to extend to the bimanual case and investigate how to learn more complex piano movements, thus being able to increase the overall difficulty of the played songs.
@@ VI Conclusion
In this work, we showcase the first proof-of-concept of how to learn to play piano with a real world robotic hand. Specifically, we employ a Sim2Real approach where we first train a policy in simulation, and then deploy it in the real world. We also introduce Hybrid Execution, a novel approach to infer with the model in the real world. Hybrid execution allows the agent to observe and act in the real world with a reduced Sim2Real gap compared to a traditional Sim2Real transfer. Hybrid execution has the most potential in complicated setups like piano-playing, where it is hard to obtain a model robust enough to be deployed traditionally using Real World Execution. Our experimental results demonstrate that our model can play several pieces successfully in the real world. Nonetheless, there is still a long way before we can achieve in robots the same level of proficiency and elegance as human pianists. Beyond the artistic raison d’être, we believe that playing piano is a relevant and challenging benchmark for dexterous contact-reach manipulation, and advances in this task can be more widely relevant for the whole manipulation community.
@@ VI-A The Reward Formulation
The $r_{\text{energy}}$ reward penalizes fast movements that need a lot of force. Further, it avoids oscillations of fingers that don’t lead to higher rewards. The energy reward also leads to a policy that behaves more fluently, which makes it more human-like.
@@
[…]
The distances are bounded between 0 and $1$ using the tolerance function of the dm_control library [13], notated as $g$. Further is $P_{\text{targets}}$, the set of the cartesian positions of all targeted keys, and $p_{\text{palm}}$ represents the cartesian position of the palm.
@@
For $k_{\text{target}}>0$ we divide the keypress reward into three cases